What mass of oxygen gas is consumed by the reaction of 5.4g of octane?

A major component of gasoline is octane (C₈H₁₈). When octane is burned in air, it chemically reacts with oxygen gas (O₂) to produce carbon dioxide (CO₂) and water (H₂O)


What mass of oxygen gas is consumed by the reaction of 5.4g of octane? 

Round your answer to 2 significant digits.
Answer #1

5.4 g of OCtane

C8H18 + 25/2 O2 --> 9 H2O + 8CO2

1 mol of Octane needs 25/2 mol of O2

MW Octane = 114 g/gmol

MW O2 = 32 g/gmol

Mol Octane = mass OCtane / MW Octane = 5.4/114 = 0.047368 mol Octane

will need 12.5x0.047368 or 0.5921 mol of O2

But O2 = 32 g/gmol

mass = 0.5921 mol O2 * 32g/gmol = 18.947 grams of O2

18.95 grams of O2